A certain force gives an acceleration
of 4 m/s 2 to a 20-kg block. What
acceleration will be imparted by the same
force to a 15-kg block?
1
Expert's answer
2020-11-25T10:45:51-0500

Let's first find the magnitude of the force that accelerates a 20-kg block at "4\ \dfrac{m}{s^2}":


"F=ma=20\ kg\cdot 4\ \dfrac{m}{s^2}=80\ N."

Finally, we can find the acceleration imparted by the same force to a 15-kg block:


"a=\dfrac{F}{m}=\dfrac{80\ N}{15\ kg}=5.33\ \dfrac{m}{s^2}."

Answer:

"a=5.33\ \dfrac{m}{s^2}."